OpenStudy (anonymous):

can someone help me with this problem: (-3x^2+5x+2)-(3x^2-2x-3) Show all work. Find the difference

OpenStudy (turingtest):

first distribute the - to the parentheses to simplify this then combine like terms can you explain where you are having trouble?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow -(x - y - z) = -x + y + z }\) Use this property, and then do operations on like terms.

OpenStudy (anonymous):

the whole problem. never done before

OpenStudy (anonymous):

show me

OpenStudy (anonymous):

(-3x^2+5x+2)-(3x^2-2x-3)=-6x^2+7x+3 -3x^2+5x+2-3x^2--2x--3 and -- = +

Parth (parthkohli):

Why? Because -(x - y - z) = -1(x - y - z) = -1x -(-y) -(-z) = -x + y + z

OpenStudy (anonymous):

-6 x^2 + 7x +2

OpenStudy (anonymous):

so 7x+5

OpenStudy (anonymous):

so schoolboy where did you go you did not help me finish

OpenStudy (anonymous):

thats the best i can do man.

OpenStudy (anonymous):

ok schoolboy, thanks anyways.

OpenStudy (turingtest):

just write the expression after distributing the - like parth said I will help you from there

OpenStudy (turingtest):

we do not just provide the whole answer when possible the aim of this site is teaching, and we feel that is best done through guiding you to find the answer yourself

OpenStudy (anonymous):

yea i know can you just explain this one to me, because i learn better step by s

OpenStudy (anonymous):

yea i know can you just explain this one to me, because i learn better step by s

OpenStudy (anonymous):

step sorry

OpenStudy (turingtest):

I am explaining: step 1) write out the expression after distributing the - sign, -(3x^2-2x-3) ^^^simplify

OpenStudy (anonymous):

9x-5x-2= 9x+-5x+-2= 4x+-2= 2x

OpenStudy (anonymous):

is this right

OpenStudy (anonymous):

please someone help me

OpenStudy (turingtest):

distributing the - gives (-3x^2+5x+2)-(3x^2-2x-3)=-3x^2+5x+2-3x^2+2x+3 now combine like terms